How do I graph y=(x+3)^2-9? Since I don't have a middle coefficient...?
Answer by wilft1(217)

solve the same way you would solve any other x,y
where did i get the 2 from? i substituted the x for 2 and then solved for y
look at the second one i did, i did the same thing except i substituted the x for 5 this time. thats how you get the coordinates, substitute one variable and then solve for the other.
y=(x+3)^2-9
y= (2+3)^2-9
y= (5)^2-9
y= 25-9
y= 14
x y
(2,14)
y=(x+3)^2-9
y=(5+3)^2-9
y=(8)^2-9
y= 64-9
y= 55
x y
(5,55)